# SWAT: Sliding Window Adversarial Training for Gradual Domain Adaptation

### v1

@@ Abstract
Domain shifts are critical issues that harm the performance of machine learning. Unsupervised Domain Adaptation (UDA) mitigates this issue but suffers when the domain shifts are steep and drastic. Gradual Domain Adaptation (GDA) alleviates this problem in a mild way by gradually adapting from the source to the target domain using multiple intermediate domains. In this paper, we propose Sliding Window Adversarial Training (SWAT) for Gradual Domain Adaptation. SWAT uses the construction of adversarial streams to connect the feature spaces of the source and target domains. In order to gradually narrow the small gap between adjacent intermediate domains, a sliding window paradigm is designed that moves along the adversarial stream. When the window moves to the end of the stream, i.e., the target domain, the domain shift is drastically reduced. Extensive experiments are conducted on public GDA benchmarks, and the results demonstrate that the proposed SWAT significantly outperforms the state-of-the-art approaches. The implementation is available at: https://anonymous.4open.science/r/SWAT-8677.
Machine Learning, ICML
@@ 1 Introduction
Conventional machine learning methods or approaches frequently presuppose that training and testing data adhere to a uniform distribution, thereby ensuring the model’s generalizability to novel data. However, this assumption is not always valid in practical applications, where domain shifts—defined as discrepancies between the source and target domains—can significantly impair model performance (Farahani et al., 2021). To illustrate, an image classifier developed using images from a controlled laboratory may have reduced accuracy when deployed on real-world images captured under varying lighting or weather conditions (Ganin & Lempitsky, 2015b; Tzeng et al., 2017).
To address this issue, Unsupervised Domain Adaptation (UDA) proposed to mitigate domain shifts by aligning feature distributions between a labeled source domain and an unlabeled target domain (Pan & Yang, 2009; Hoffman et al., 2018). Nevertheless, Kang et al. (2019); Tang & Jia (2020); Yang et al. (2020) have revealed that when the domain gaps are large, direct alignment of the source and target domains can negatively impact the model and cause the negative transfer. Direct feature-level alignment between source and target domains cause geometric distortion from enforcing rigid distribution matching (e.g., MMD), where forcing alignment of non-overlapping support sets amplifies classifier boundary errors (Zhao et al., 2019). Additionally, static alignment blindness ignores the latent domain evolution trajectory, leading to suboptimal adaptation paths; and discriminability erosion where excessive invariance learning causes loss of category-aware structures (Liang et al., 2020).
Figure 1: Overview of gradual domain adaptation. The classifiers perform accurately in classifying the sample points in the current domain and the neighboring domains, but is fail when classifying the samples from the Source directly to the Target domain.
Gradual Domain Adaptation (GDA)(Kumar et al., 2020) is proposed to alleviate this problem in a mild way by gradually adapting from the source to the target domain using multiple intermediate domains, as shown in Fig. 1. Recently, GDA advances include theoretical analyses of error propagation in self-training (Wang et al., 2022), optimal transport-guided intermediate domain generation via Wasserstein geodesics (GOAT) (He et al., 2023), and normalizing flow-based feature alignment (Sagawa & Hino, 2022). Nevertheless, current GDA methods face two fundamental bottlenecks: (1) Existing GDA frameworks rely on discretized intermediate domains (e.g., fixed-step interpolations), leading to quadratic error accumulation (Kumar et al., 2020; He et al., 2023) due to noise propagation in pseudo-labeling across stages. Rigid phase divisions further disrupt feature manifold continuity, causing semantic distortion and degraded transfer fidelity. (2) Manually designed intermediate domains (He et al., 2023) often fail to align with real-world domain shift dynamics. Uniform interpolation assumptions (e.g., linear paths) introduce spatial redundancy (ineffective transition regions) and density imbalance (over-dispersion or mode collapse), destabilizing adaptation trajectories (Xiao et al., 2024).
We present a novel framework for gradual domain adaptation, dubbed Sliding Window Adversarial Training (SWAT). This framework progressively transfers knowledge from the source domain to the target domain through adversarially transported intermediate representations. SWAT constructs a feature transport flow by iteratively aligning the model to a sequence of dynamically adjusted intermediate domains. This flow is optimized by curriculum-guided window sliding, enabling smooth transitions. Specifically, our contributions are as follows:
@@
    Sliding Window Adversarial Training: We incorporate adversarial learning with a curriculum sliding window mechanism. Instead of fixed interpolations, SWAT employs a sliding window that adaptively focuses on regions along the source-to-target path.
@@
    Adversarial Flow Matching: We propose a bidirectional adversarial framework unifying temporal flow alignment and feature matching. The flow generator enforces domain continuity through sliced Wasserstein optimization across evolving domains, while the discriminator progressively filters out source-specific features through adversarial attenuation. This co-evolutionary optimization achieves simultaneous domain invariance and target discriminability.
@@
    Experimental Superiority: Experiments on Rotated MNIST (96.7% vs. 86.4% SOTA) and Portraits (87.4% vs. 86.16% SOTA) demonstrate superiority, with ablation studies showing 36.0% error reduction in extreme shifts (Fig. LABEL:fig:abl-subfig2).
@@ 2 Related Work
Unsupervised Domain Adaptation (UDA) aims to align feature distributions between labeled source domains and unlabeled target domains to mitigate domain shifts. Traditional approaches leverage statistical measures such as Maximum Mean Discrepancy (MMD) (Chen et al., 2020) to enforce domain invariance, while facing fundamental limitations under severe distribution divergence, despite their effectiveness for moderate domain gaps. Rigid MMD-based alignment may forcibly align non-overlapping supports, distorting classifier boundaries (Zhao et al., 2019), while direct source-target alignment risks erasing category-discriminative structures—a phenomenon termed negative transfer (Tang & Jia, 2020; Yang et al., 2020). Such limitations arise from static alignment strategies that ignore the geometric continuity of latent domain trajectories. Unlike static UDA frameworks, ours SWAT dynamically adjusts alignment granularity dynamically via a Wasserstein-based curriculum.
@@
Adversarial Domain Adaptation frameworks, including DANN (Ganin & Lempitsky, 2015a) and CDAN (Long et al., 2018), have revolutionized alignment through adversarial training. These methods employ gradient reversal layers or conditional adversarial networks to learn domain-invariant representations. However, these methods enforce fixed pairwise alignment between source and target domains, leading to mode collapse when domain supports are disjoint (Zhao et al., 2019) or under large distribution gaps due to gradient competition (Pezeshki et al., 2021). Recent advances, such as spectral regularization (Pezeshki et al., 2021), partially alleviate these issues but retain the rigidity of discrete alignment steps. In contrast, our SWAT redefines domain adaptation as a continuous manifold transport process. By constructing intermediate domains along a feature transport flow, SWAT avoids abrupt transitions and assimilates novel target modes progressively—a critical failure point for conventional UDA and adversarial methods alike.
Gradual Domain Adaptation (GDA) addresses scenarios where data shifts gradually, decomposing the overall shift into smaller steps via intermediate domains (Farshchian et al., 2018; Kumar et al., 2020). While existing methods leverage self-training (Xie et al., 2020), gradient flow-based geodesic paths (Zhuang et al., 2024), style-transfer interpolation (Marsden et al., 2024), or optimal transport (He et al., 2023), they often suffer from catastrophic forgetting of source knowledge during incremental adaptation. Our SWAT framework uniquely preserves source-acquired information through adversarial flow calibration, where Wasserstein-guided intermediate domains progressively integrate target features.
Domain Flow Generation Model (DLOW) proposes to synthesize intermediate domains between source and target domains through adversarial learning (Gong et al., 2021). The core idea involves training bidirectional generators that progressively transform data distributions from the source to target domain (and vice versa) under a parametric control variable. By incorporating adversarial discriminators and cycle-consistency constraints inspired by CycleGAN (Zhu et al., 2017). This framework demonstrates the feasibility of continuous domain interpolation. Our method uses an adversarial encoder to bridge the distance between domains, while using a Wasserstein distance constraint to transfer paths. Trajectory learning in latent space can maintain consistency across domains.
@@ 3 Problem Setup
In the gradually domain setting(Kumar et al., 2020), given a sequence of domains $\{P_{t}\}_{t=0}^{n}$ with gradually shifting distributions, where $P_{0}$ is the labeled source domain and $P_{n}$ the unlabeled target domain, GDA aims to learn a hypothesis $h:\mathcal{X}\to\mathcal{Y}$ that minimizes target risk $\epsilon_{n}(h)$, under two core assumptions (Kumar et al., 2020; Long et al., 2015): Bounded Successive Divergence and Conditional Invariance.
$\text{W}_{1}(P_{t},P_{t+1})\leq\Delta,\quad\forall t\in\{0,...,n-1\}$
$P_{t}(Y|X)=P_{t+1}(Y|X),\quad\forall t\in\{0,...,n-1\}$
@@
The goal of pretraining in the source domain is to learn a model $C:\mathcal{X}\to\mathcal{Y}$ that maps input features $x$ from the training data set $\mathcal{D}=\{(x,y)\}$ to their corresponding labels $y$. Considering the loss function $l$, the classifier benefit on $\mathcal{D}_{t}$ is denoted by $C$, defined as:
@@
The goal of gradual domain adaptation is to learn a classifier $C$ that generalizes well to the target domain $\mathcal{D}_{n}$ by progressively transferring knowledge from the labeled source domain $\mathcal{D}_{0}$ and a series of unlabeled intermediate domains $\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{n-1}$. The adaptation process involves multi-step pseudo-labeling and self-training, where the model $C_{0}$ is trained on the source domain and then adapted to the intermediate domains by the following self-training procedure $\text{ST}(C_{t},\mathcal{D}_{t})$:
@@
Figure 2: The SWAT framework incrementally aligns feature distributions across domains via adversarial training. Blue dashed lines denote bidirectional adversarial interactions (e.g., domain discriminators and task classifiers) that enforce domain invariance while preserving target discriminability. This dual learning process enables gradual adaptation by balancing cross-domain alignment and task-specific performance.
@@ 4 Methodology
4.1 Method Overview
The proposed Sliding Window Adversarial Training (SWAT) models gradual domain adaptation as a continuous feature flow matching process guided by adversarial learning. SWAT integrates three core mechanisms. First, As depicted in Fig. 3, the framework establishes a time-conditional probability path $\{p_{t}(\mathbf{h})\}_{t\in[0,n]}$ over latent space $\mathcal{H}\subseteq\mathbb{R}^{z}$, where the encoder $f_{t}$ and classifier $g_{t}$ evolve smoothly from source ($f_{0},g_{0}$) to target ($f_{n},g_{n}$). Second, a bidirectional adversarial architecture (eq. 5) employs a forward generator $G_{m}$ to map source features $\mathcal{H}_{0}$ to intermediate domains via Wasserstein-optimal transport, and a reverse generator $G_{s}$ to project features back to $\mathcal{H}_{0}$ for cycle consistency (Eq. 6), which ensure distributional alignment. Third, discriminator $D$ and encoder $f$ form an adversarial architecture to encourages domain-invariant features, enabling SWAT to balance stability and flexibility during domain shifts.
4.2 Continuous Feature Flow Matching
As shown in Fig. 3, $\mathcal{H}\subseteq\mathbb{R}^{z}$ is a $z$-dimensional space feature, $g\circ f$ denotes the model, where $f\in\mathcal{X}\to\mathcal{H}$ is the encoder which maps the input data into the latent feature space $\mathcal{H}$, while $g\in\mathcal{H}\to\mathcal{Y}$ serves as the classifier, which predicts the labels based on the encoded features. Specifically, when traversing the source, intermediate, and target domains, the encoders are represented as $f_{0},f_{1},\ldots,f_{n}$. Similarly, the classifiers can be represented as $g_{0},g_{1},\ldots,g_{n}$, where $f_{t}$ and $g_{t}$ denote the model that continuously evolves across domains.
Figure 3: Illustration of sliding window mechanism. The mid-ground represents the sample space with distinct domains $\mathcal{D}_{i}$ (where $i=0,\ldots,4$) positioned across it. The feature flows, depicted as interconnected cards in the middle, are characterized by the encoder and reside within a contiguous representation space. Above, the linear progression of features in the feature space $\mathcal{H}_{i}$ is shown.
The feature domains produced after encoding by $f$ are represented as $\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{n}$. $f$, $g$, and $\mathcal{H}$ encapsulate a series of continuous transformations and domains that evolve progressively to facilitate adaptation from the source domain to the target domain.
4.3 Sliding Window Adversarial Training (SWAT)
The proposed sliding window adversarial training is visualized in Fig. 2. Unlike conventional discrete domain transfers, SWAT enables continuous feature transfering along the domain stream $\mathcal{H}^{z}$ ($z\in[0,n]$) through sliding window. The core idea lies in the parametric sliding window defined by $p\in[0,1]$, which controls the relative position between $\mathcal{H}^{z}$ and its neighboring critical domains $\mathcal{H}_{l}$, $\mathcal{H}_{r}$ ($r=l+1$). This creates adaptive transmission paths $\mathcal{H}_{0}\leftrightarrow\mathcal{H}^{(l+p)}$ where the target domain smoothly shifts with $p$.
Here, $l\in\{0,1,\dots,n-1\}$ denotes the index of the left neighboring domain, and the parameter $p$ controls the distance from the source domain to the left critical domain. In particular, when $p=0$, the transfer occurs to the left domain $\mathcal{H}_{l}$; when $p=1$, the transfer occurs to the right domain $\mathcal{H}_{r}$. The expression $\mathcal{H}^{(l+p)}$ refers to a domain located between $\mathcal{H}_{l}$ and $\mathcal{H}_{r}$, with $r=l+1$ representing the index of the neighboring right domain.
@@ 4.3 Sliding Window Adversarial Training (SWAT)
In our cross-domain translation model, we define $G_{m}$ as the transformation function that maps a sample $h$ from the source domain $\mathcal{H}_{s}$ to a target domain within the domain stream $\mathcal{H}^{z}$, where $z\in[0,n]$ indicates the position of the target domain within the stream. Conversely, $G_{s}$ denotes the reverse transformation, mapping features from any domain in the stream $\mathcal{H}^{z}$ back to the source domain $\mathcal{H}_{s}$. Thus, our SWAT model can be expressed as the bidirectional transformations:$G_{m}:\mathcal{H}_{s}\to\mathcal{H}^{z}$ and $G_{s}:\mathcal{H}^{z}\to\mathcal{H}_{s}$.
@@
As in the DLOW model described in Section 2, we employ a Generative Adversarial Network (GAN) (Goodfellow et al., 2014) for our cross-domain transport model. Instead of using a standard GAN, we opt for the Wasserstein GAN (WGAN) (Arjovsky et al., 2017), as the Wasserstein distance (W-distance) provides a more effective measure of the distance between domains, generates higher-quality target domains $\mathcal{H}^{z}$, and is easier to train.
The objective function for the adversarial training module is defined as:
$\min_{D}\max_{G}\mathbb{E}_{\begin{subarray}{c}\hat{h}\sim\mathbb{P}_{g}\\ h\sim\mathbb{P}_{r}\end{subarray}}\left[D(\hat{h})-D(h)\right]+\mathcal{R}.$ (3)
@@
Where $\hat{h}$ represents a sample generated by the generator $G$, which approximates the target domain distribution $\mathbb{P}_{g}$. $h$ is a sample from the real data distribution $\mathbb{P}_{r}$, corresponding to actual data from the target domain. $D$ denotes the discriminator of the corresponding domain, and different domains have different discriminators. $D$ is trained to distinguish between real samples $h$ and fake samples $\hat{h}$, with the goal of minimizing the Wasserstein distance between the two distributions. And $\mathcal{R}$ represents the regularization term proposed by Gulrajani et al. (2017). Specifically, the regularization term is given by:
$\mathcal{R}=\mathbb{E}_{\tilde{h}\sim\mathbb{P}_{\tilde{h}}}\left[\lambda\left(||\nabla_{\tilde{h}}D(\tilde{h})||_{2}-1\right)^{2}\right].$ (4)
@@
Here, $\tilde{h}$ denotes a random linear interpolation of points from $\bar{h}$ and $h$ samples, and $\lambda$ is a hyperparameter controlling the strength of the regularization.
Adversarial Domain Alignment
@@
To promote bidirectional feature alignment between the source domain $\mathcal{H}_{0}$ and critical domains, we design adversarial losses based on the minimax objective $V(\mathbb{P}_{g},\mathbb{P}_{r})$ defined in Eq. (3). For the left critical domain $\mathcal{H}_{l}$, the adversarial loss enforces cross-domain distribution matching through dual mapping paths:
$\mathcal{L}^{l}_{\text{adv}}=\underbrace{V\left(G_{m}(\mathcal{H}_{0}),\mathcal{H}_{l}\right)}_{\text{Forward mapping}}+\underbrace{V\left(G_{s}(\mathcal{H}_{l}),\mathcal{H}_{0}\right)}_{\text{Backward reconstruction}},$ (5)
@@
To prevent mode collapse and maintain content integrity during cross-domain translation, we employ cycle-consistent constraints inspired by CycleGAN. This ensures that features cyclically transformed through $\mathcal{H}_{0}\rightarrow\mathcal{H}_{l}\rightarrow\mathcal{H}_{0}$ preserve semantic equivalence:
$\begin{gathered}\mathcal{L}^{l}_{\text{cycle}}=\mathbb{E}_{h\sim\mathcal{H}_{0}}\left[\|G_{s}(G_{m}(h))-h\|_{2}\right]\\ +\mathbb{E}_{h\sim\mathcal{H}_{l}}\left[\|G_{m}(G_{s}(h))-h\|_{2}\right],\end{gathered}$ (6)
The bidirectional reconstruction constraints enforce invertible transformations while penalizing semantic distortions, particularly crucial for preserving task-relevant features in critical domains.
Figure 4: Illustration of the domain flow with multiple potential paths connecting the domains. The red dot represent the generated intermediate domain $\mathcal{H}^{(l+p)}$, which lies along the shortest domain flow between $\mathcal{H}_{l}$ and $\mathcal{H}_{r}$. The placement of $\mathcal{H}^{(l+p)}$ are determined by $p$.
4.4 The Overall Objective for Gradual Domain Adaptation
Figure 5: This architecture introduces a bidirectional transport framework for domain adaptation, constructing an intermediate domain $\mathcal{H}^{(l+p)}$ through coupled generators $G_{s}$ (upward flow) and $G_{m}$ (downward flow). Key components include: 1) a residual-connected fully connected network $G$, 2) a three-convolutional-layer encoder, and 3) dual adversarial modules ($D$, $g$) with three-layer fully connected structures. Pseudo-source domains $\mathcal{H}^{\prime}_{l}/\mathcal{H}^{\prime}_{r}$ are synthesized from source features $\mathcal{H}_{l}/\mathcal{H}_{r}$, governed by adaptive loss weights $(1-p)$ and $p$ for left/right branches. Sliding window adaptation mechanism (lower-left arrow) enables $\mathcal{D}_{l}$ and $\mathcal{D}_{r}$ scrolling forward step by step.
Similar to previous GDA mehtods, we also optimize the self-training loss as follows:

### v2

@@ Abstract
Domain shifts are critical issues that harm the performance of machine learning. Unsupervised Domain Adaptation (UDA) mitigates this issue but suffers when the domain shifts are steep and drastic. Gradual Domain Adaptation (GDA) alleviates this problem in a mild way by gradually adapting from the source to the target domain using multiple intermediate domains. In this paper, we propose Sliding Window Adversarial Training (SWAT) for GDA. SWAT first formulates adversarial streams to connect the feature spaces of the source and target domains. Then, a sliding window paradigm is designed that moves along the adversarial stream to gradually narrow the small gap between adjacent intermediate domains. When the window moves to the end of the stream, i.e., the target domain, the domain shift is explicitly reduced. Extensive experiments on six GDA benchmarks demonstrate the significant effectiveness of SWAT, especially 6.1% improvement on Rotated MNIST and 4.1% advantage on CIFAR-100C over the previous methods.
@@ 1 Introduction
Traditional machine learning assumes identical training-test data distributions, yet real-world domain shifts often break this assumption and degrade model performance [1]. Unsupervised Domain Adaptation (UDA) is proposed to mitigate domain shifts by aligning feature distributions between a labeled source domain and an unlabeled target domain [2, 3]. Nevertheless, existing works [4, 5, 6] have revealed that when the domain gaps are large, directly aligning two domains not only fails to reduce the domain gaps, but even causes the negative transfer [2].
Gradual Domain Adaptation (GDA)[7] is proposed to alleviate this problem in a mild way by gradually adapting from the source to the target domain using multiple intermediate domains, as shown in Fig. 1. This paper addresses the GDA problem through adversarial training. Adversarial training has been widely used in UDA and achieved impressive performance. This training paradigm, however, faces two challenges when applying in GDA. On the one hand, previous adversarial training methods (e.g., DANN[8]) globally align two distributions through the game between generator and discriminator. This global matching cannot handle the continuous intermediate domains in GDA [9, 10]. As a result, the GDA problem degrades to the more difficult UDA problem.
On the other hand, the steep gradient of the adversarial training for large domain shifts will cause discontinuities and unsmooth problems in the manifold space [11, 10, 12]. As machine learning methods rely on the continuous and smooth manifold hypothesis to avoid abrupt changes in decision boundaries, this discontinuity and unsmoothness will cause error accumulation [7, 13, 14].
Towards the smooth and stable distribution matching, we propose the sliding window mechanism for adversarial learning. As a new training paradigm, the sliding window mechanism emerges three advantages over the traditional adversarial training: (1) Locality: The sliding window mechanism avoids the globally alignment by localizing the adversarial training range, i.e., it decomposes the continuous domain flow into multiple windows, and perform adversarial training in each window to gradually refine the alignment. The generator only focuses on the distribution differences in the current window, which reduces the complexity of the adversarial training. (2) Dynamic: The window is gradually shifted from the source domain to the target with the training process, and the update frequency of generator parameters is synchronized with the speed of domain alignment, avoiding the error accumulation caused by large optimization step in traditional adversarial training. (3) Continuity: The sliding window continuously slides on the domain stream $H_{z}$, and the continuous change of parameter $p$ from 0 to 1 gradually turns the optimization focus from the left domain $H_{l}$ to the right domain $H_{r}$, avoiding the discrete switching in the multi-stage training.
Incorporated the sliding window mechanism, we present Sliding Window Adversarial Training (SWAT) method for GDA. Specifically, SWAT first formulates a bidirectional adversarial flow. This flow is optimized by a curriculum-guided sliding window, which finely controls the transition step between the source and the target domains, avoiding quadratic error accumulation caused by large transfer steps of the existing self-training strategy [7]. In the adversarial training phase, the flow generator enforces domain continuity through sliced Wasserstein optimization across evolving domains, while the discriminator progressively filters out source-specific features through adversarial training. This co-evolutionary optimization achieves simultaneous domain invariance and target discriminability[14]. The contributions are summarized as follows:
@@
    We propose a sliding window mechanism to improve the adversarial training, which splits large domain shifts into multiple micro transfers through local, dynamic and continuous feature alignment, enabling fine-grained distribution matching.
@@
    We present the Sliding Window Adversarial Training (SWAT) method for GDA. SWAT can adaptively align localized domain regions, mitigating error accumulation and enabling smooth and robust knowledge transfer.
@@
    Experiments on Rotated MNIST (96.7% vs. 90.6% SOTA), Portraits (87.4% vs. 86.16% SOTA) and CIFAR-100C (24.8% vs. 28.9%) demonstrate the effectiveness of SWAT.
Figure 1: Comparison of UDA and GDA. Left (UDA): A single alignment maps source features directly onto the target domain. Right (GDA): Adaptation proceeds through a sequence of intermediate domains that smoothly adapt across domains, reducing abrupt distribution shifts.
@@ 2 Related Work
Unsupervised Domain Adaptation (UDA) aims to mitigate domain shifts by aligning feature distributions between labeled source and unlabeled target domains. Traditional approaches leverage statistical measures like Maximum Mean Discrepancy (MMD) [15] to enforce domain invariance, but face limitations under severe distribution divergence: rigid MMD-based alignment risks distorting classifier boundaries by forcibly aligning non-overlapping supports [16], while direct source-target alignment may erase discriminative structures, causing negative transfer [5, 6]. Adversarial methods like DANN [17] and CDAN [18] advanced alignment via adversarial training but enforce fixed pairwise alignment, leading to mode collapse under disjoint supports [16] or gradient competition under large gaps [19]. While spectral regularization [19] partially alleviates these issues, it retains rigid alignment steps.
Gradual Domain Adaptation (GDA) addresses scenarios where domain shifts occur incrementally, decomposing the overall distribution gap into smaller, more manageable steps through intermediate domains [20, 7]. Existing methods employ diverse strategies to model these transitions: self-training leverages pseudo-labeling to bootstrap target predictions [21], gradient flow-based geodesic paths enforce smooth transitions via Riemannian manifolds [22], style-transfer interpolation synthesizes intermediate domains through low-level feature mixing [23], and optimal transport (OT) aligns domain distributions using Wasserstein distances [13]. While alignment alone may rigidly match marginal distributions at the expense of discriminative structures. These issues are exacerbated in multi-step adaptation, where imperfectly aligned intermediates compound errors, leading to irreversible distortion of decision boundaries. Our SWAT framework uniquely preserves source-acquired information through bidirectional alignment, balancing between stability and plasticity,
@@
Adversarial Domain Adaptation frameworks, including DANN [17] and CDAN [18], have revolutionized alignment through adversarial training. These methods employ gradient reversal layers or conditional adversarial networks to learn domain-invariant representations. However, these methods enforce fixed pairwise alignment between source and target domains, leading to mode collapse when domain supports are disjoint [16] or under large distribution gaps due to gradient competition [19]. Recent advances, such as spectral regularization [19], partially alleviate these issues but retain the rigidity of discrete alignment steps. In contrast, our SWAT redefines domain adaptation as a continuous manifold transport process. By constructing intermediate domains along a feature transport flow, SWAT avoids abrupt transitions and assimilates novel target modes progressively, i.e., a critical failure point for conventional UDA and adversarial methods alike.
@@ 3 Problem Setup
In the gradually domain setting[7], given a sequence of domains $\{P_{t}\}_{t=0}^{n}$ with gradually shifting distributions, where $P_{0}$ is the labeled source domain and $P_{n}$ the unlabeled target domain, GDA aims to learn a hypothesis $h:\mathcal{X}\to\mathcal{Y}$ that minimizes target risk $\epsilon_{n}(h)$, under two core assumptions [7, 24]: (1) the distributional shift between consecutive domains is limited, known as Bounded Successive Divergence, and (2) the conditional distribution of labels given inputs remains unchanged across domains, referred to as Conditional Invariance.
$\text{W}_{1}(P_{t},P_{t+1})\leq\Delta,\quad P_{t}(Y|X)=P_{t+1}(Y|X),\quad\forall t\in\{0,.,n-1\},$ (1)
@@
The goal of pretraining in the source domain is to learn a model $C:\mathcal{X}\to\mathcal{Y}$ that maps input features $x$ from the training data set $\mathcal{D}=\{(x,y)\}$ to their corresponding labels $y$. Considering the loss function $l$, the classifier optimized on $\mathcal{D}_{t}$ is denoted by $C$, defined as:
@@
Gradual domain adaptation aims to train a model $C$ that effectively generalizes to the target domain $\mathcal{D}_{n}$ by incrementally transferring knowledge from the labeled source domain $\mathcal{D}_{0}$ through a sequence of unlabeled intermediate domains $\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{n-1}$. The adaptation process involves multi-step pseudo-labeling and self-training, where the model $C_{0}$ is trained on the source domain and then adapted to the intermediate domains by the following self-training procedure $\text{ST}(C_{t},\mathcal{D}_{t})$:
@@ 4 Methodology
(a) Feature Flow Matching
(b) Cross-domain Adversarial Training
Figure 2: (a) Illustration of sliding window mechanism, where overlapping feature spaces facilitate smooth domain transitions along the domain sequence. (b) Incremental domain alignment with the SWAT framework using adversarial training to preserve task performance and encourage feature consistency across domains.
The proposed sliding window adversarial training models gradual domain adaptation as a continuous transition of feature distributions guided by adversarial learning. The core components are detailed in the following.
Continuous Feature Flow
As illustrated in Fig. 2(a), SWAT first defines a continuous sequence of feature distributions:
$\left\{\mathcal{H}_{t}\right\}_{t\in[0,n]},\quad\mathcal{H}_{t}\;=\;p_{t}(\mathbf{h}),$ (4)
over the latent space $\mathcal{H}\subseteq\mathbb{R}^{z}$, where each $\mathcal{H}^{i}$ is the feature manifold at adaptation step $i$. Here $f_{t}:\mathcal{X}\to\mathcal{H}$ and $g_{t}:\mathcal{H}\to\mathcal{Y}$ denote the encoder and classifier at step $t$, so that the overall model $g_{t}\circ f_{t}$ smoothly evolves from $(f_{0},g_{0})$ on the source domain to $(f_{n},g_{n})$ on the target.
Traditional GDA methods treat each pair $(f_{i},g_{i})$ as an independent stage in the adaptation process, where $(f_{i+1},g_{i+1})$ is trained only after $(f_{i},g_{i})$ has converged. This leads to a sequence of discrete transitions between domains. In contrast, SWAT learns a continuous sequance of models $\{(f^{z},g^{z})\}_{z\in[0,n]}$ and aligns features along the entire path $\mathcal{H}^{z}$. This continuous flow matching avoids abrupt transitions and enables fine‐grained adaptation at every intermediate point.
Unlike the discrete adaptation process in previous GDA methods, SWAT enables continuous feature transferring along the domain stream $\mathcal{H}^{z}$ ($z\in[0,n]$) through a sliding window, as illustrated in Fig. 2(b).
Sliding Window Mechanism
At any step $l\in\{0,\dots,n-1\}$, sliding window is the pair of adjacent domains $\{\mathcal{H}_{l},\mathcal{H}_{r}\}$ with $r=l+1$. The scalar parameter $p\in[0,1]$ controls where within that window we align:
$\mathcal{H}^{(l+p)}\;=\;(1-p)\,\mathcal{H}_{l}\;+\;p\,\mathcal{H}_{r},$ (5)
where $\mathcal{H}^{(l+p)}$ refers to a domain located between $\mathcal{H}_{l}$ and $\mathcal{H}_{r}$. Here, $\mathcal{H}_{l}$ and $\mathcal{H}_{r}$ refer to the left and right critical domains. When $p$ reaches 1, the window “slides” one step to the right (i.e. $l\leftarrow l+1$ and $p$ resets toward 0), so that the next window becomes $\{\mathcal{H}_{l+1},\mathcal{H}_{l+2}\}$. As $p$ varies from 0 to 1 and then triggers a slide, SWAT sweeps continuously through the entire domain stream. We then formalize the sliding-path alignment as:
$\mathcal{H}_{0}\leftrightarrow\mathcal{H}^{(l+p)},\quad l\in\{0,1,\dots,n-1\},\quad p\in[0,1],$ (6)
where both $l$ and $p$ are parameters controlling smooth transitions across domains. This formulation enables fine-grained domain alignment through continuously shifting intermediate representations.
Bidirectional Flow Matching
@@
Building upon the sliding window mechanism, we further incorporate it with the adversarial training for smooth flow matching. Specifically, we define $G_{m}$ as the transformation function that maps a representation $h$ from the source domain representation $\mathcal{H}_{s}$ to a target domain within the domain stream $\mathcal{H}^{z}$, where $z\in[0,n]$ indicates the position of the target domain within the stream. Conversely, $G_{s}$ denotes the reverse transformation, mapping features from any domain in the stream $\mathcal{H}^{z}$ back to the source domain $\mathcal{H}_{s}$. Thus, our SWAT model can be expressed as the bidirectional transformations:
$G_{m}:\mathcal{H}_{s}\to\mathcal{H}^{z},\quad G_{s}:\mathcal{H}^{z}\to\mathcal{H}_{s}.$ (7)
@@
We employ the Wasserstein GAN (WGAN) [25] to train the SWAT model, as the Wasserstein distance provides a more effective measure of the distance between domains, generates higher-quality target domains $\mathcal{H}^{z}$, and is easier to train. The objective function for the adversarial training module is defined as:
$\min_{D}\max_{G}V(\mathbb{P}_{g},\mathbb{P}_{r})=\min_{D}\max_{G}\mathbb{E}_{\begin{subarray}{c}\hat{h}\sim\mathbb{P}_{g}\\ h\sim\mathbb{P}_{r}\end{subarray}}\left[D(\hat{h})-D(h)\right]+\mathcal{R},$ (8)
@@
where $\hat{h}$ represents a representation generated by the generator $G$, which approximates the target domain distribution $\mathbb{P}_{g}$. $h$ is a representation from the real data distribution $\mathbb{P}_{r}$, corresponding to actual data from the target domain. $D$ denotes the discriminator of the corresponding domain, and different domains have different discriminators. $\mathcal{R}$ represents the regularization term proposed by Gulrajani et al. [26]:
[…]
@@
where $\tilde{h}$ denotes a random linear interpolation of points from $\hat{h}$ and $h$ representations, and $\lambda$ is a hyperparameter controlling the strength of the regularization.
@@
To facilitate bidirectional feature alignment between the source domain $\mathcal{H}_{0}$ and the critical domains, we formulate bidirectional flow matching based on the minimax objective $V(\mathbb{P}_{g},\mathbb{P}_{r})$ defined in Eq. (8). Without loss of generality, taking the left critical domain $H_{l}$ as an example, the adversarial loss enforces cross-domain distribution matching through dual mapping paths:
$\mathcal{L}^{l}_{\text{adv}}=V\left(G_{m}(\mathcal{H}_{0}),\mathcal{H}_{l}\right)+V\left(G_{s}(\mathcal{H}_{l}),\mathcal{H}_{0}\right),$ (10)
@@
To prevent mode collapse and maintain content integrity during adaptation, we employ cycle-consistent constraints inspired by CycleGAN. This ensures that features cyclically transformed through $\mathcal{H}_{0}\rightarrow\mathcal{H}_{l}\rightarrow\mathcal{H}_{0}$ to preserve semantic consistency:
$\begin{gathered}\mathcal{L}^{l}_{\text{cycle}}=\mathbb{E}_{h\sim\mathcal{H}_{0}}\left[\|G_{s}(G_{m}(h))-h\|_{2}\right]+\mathbb{E}_{h\sim\mathcal{H}_{l}}\left[\|G_{m}(G_{s}(h))-h\|_{2}\right].\end{gathered}$ (11)
The bidirectional reconstruction regularizations enforce invertible transformations while penalizing semantic distortions, particularly crucial for preserving task-relevant features in critical domains.
4.1 The Overall Objective
Similar to previous GDA methods, we also optimize the self-training loss as follows: